OpenStudy (anonymous):

Please Help I need Help! Not just mentally but with this math problem!! AHH I am going to go insane. I have had 3 different people help me with this problem and the answers that I got are not the ones that I have to choose from. An object is moving along a horizontal line on the interval 1 ≤ t ≤ 3, where t is measured in seconds. The object has an initial position x(1) = 3 meters and a velocity of v(t) = 1.5t meters per second. Find the object’s position when t = 3 seconds

OpenStudy (anonymous):

Someone told me earlier that it was (1.5)(3)=4.5+3 = 7.5 but that is not the correct answer.

OpenStudy (anonymous):

Well v(t) = 1.5t Therefore the position x(t) can be obtained by integrating v(t). x(t) = (1.5/2)t^2 + c x(t) = 0.75t^2 + c But since we have the coordinates (1,3) 3 = 0.75(1)^2 + c c = 2.25 x(t) = 0.75t^2 + 2.25 x(3) = 0.75*3^2+ 2.25 x(3) = 9

OpenStudy (anonymous):

Thank you so much! I knew I had to integrate something, I just was not sure what.

OpenStudy (anonymous):

Yeah, 1.5(3) only applies for constant velocity. In physics based questions, the velocity is the derivative of the position And the acceleration is the derivative of the velocity